# Tension-induced giant actuation in unstructured elastic sheets

### v1

@@ Appendix A Experimental methods
The stretching equipment was set up by two linear-stage motors with $300\mathrm{~{}mm}$ range of motion fixed onto a stand, and two specimen clamp fixtures each attached to the slider module on a linear stage (Fig. A1a). Both the specimen fixtures and the stand were 3D printed (Raise3D FDM Printer) using polylactic acid (PLA). The specimen fixtures were custom designed to fit the modules on the linear stage and the clamp surface were patterned with teeth-like structures to enhance grip (Fig. A1a, inset). The specimens (Fig. A1b) were made of laser-cut commercial polyvinyl chloride (PVC) sheets. Elastic modulus of the sheets were determined from tensile tests of rectangular strips with $W_{\mathrm{load}}=W$. Averaging five samples for each thickness, we obtain the Young’s modulus of PVC sheets with thickness $0.04,0.10,0.15\mathrm{~{}mm}$ to be $3.84,3.50,1.96\mathrm{~{}GPa}$, respectively (Fig. A1c). The sheets used in the experiment have extended ’tabs’ which are regions for clamping with width of $W_{load}=10\mathrm{~{}mm}$ and $L_{load}=30\mathrm{~{}mm}$. The connection between the clamping regions and the main part of the sheets was designed with rounded corners of radius $1\mathrm{~{}mm}$ to prevent failure due to stress concentration. Tensile strains in the elastic sheets were measured by strain gauges (Tokyo Measuring Instruments Laboratory, GFLA-6-350-70) with gauge length of $6\mathrm{~{}mm}$, resistance of $350~{}\Omega$, and resolution of $1\mathrm{~{}\mu\epsilon}$. The strain gauges were attached to the center of the specimen along the loading direction using an adhesive, and the gauge leads were connected to electrical wires that input to the strain data logger (Tokyo Measuring Instruments Laboratory, TDS-303). When clamping specimens between fixtures, vertical alignment is ensured for symmetric loading. Cannula needles with an outer diameter of $700\mathrm{~{}\mu m}$ were adhered to the $y$-middle of the sheet near the edges as an indicator in the measurement of folding angles. Photographs were taken from top-view of the specimens during the loading process (Figs. 3e and A1d), from which folding angles were measured using MATLAB.
@@
Extended Data Fig. A1: Experimental setups. a, Tensile-stretching equipment consisting of linear-stage motors, 3D-printed specimen clamp fixtures and a stand (all labelled on the figure). b, Testing specimen, showing the location of the strain gauge, cannula and where the sheet is clamped (see labels). c, Stress-strain curves of the PVC sheet with fully clamped edge. d, Composite image showing the evolution of the folding angle, $\alpha$, with increasing strain.
Appendix B Finite element method simulation
@@
Numerical simulation was conducted using the commercial finite element analysis software ABAQUS/2023. The model was constructed in 3D space using 2D shell sections and S4R elements. Linear elastic constitutive relation was applied with Young’s modulus $E$=3000 MPa and Poisson’s ratio $\nu$=0.3. Displacement loading $U_{2}$ was applied to the bottom edge of the lower clamping region, while the top edge of the upper clamping region was fixed in all degrees of freedom. Two analysis steps were conducted in sequence: first, a small out-of-plane displacement to the midline at $x=0$ of a magnitude of $\sim$ 1% of shell thickness, creating initial perturbation to the planar shell; second, the static analysis step was performed for the process of tensile-stretching induced folding of the elastic sheet.
@@
Extended Data Fig. A2: FE simulation results. a, Von-Mises stress distributions in stretched sheets in Fig. 2a insets (i-iii). b-c, Out-of-plane deflections $\zeta$ normalized by thickness $t$ (b) and transverse stress $\sigma_{xx}$ (c) in stretched sheets $L/W=2,t/W=0.005$ with partially clamped edges $W_{\mathrm{load}}/W=0.2$ (left) and fully clamped edges $W_{\mathrm{load}}/W=1$ (right). Contours shown in b-c are at the critical buckling strain for the partially clamped case, i.e., $\bar{\epsilon}_{c}\approx 1.2\times 10^{-3}$.
@@
Extended Data Fig. A3: The scaled folding angle exhibits an intersection with the $x$-axis prior to the linear region, from where we determine the critical buckling strains $\bar{\epsilon}$ in Fig. 4b. They have varying $L/W$, $W_{\mathrm{load}}/W$, following the legends in Fig. 2b, and $t/W$ indicated by the color bar.
Appendix C Comparison of different strains in experiments
@@
These numerical results allow us to compare the various types of strain that are measured experimentally.
@@
Firstly, our experiments measured the local strain, $\epsilon_{\mathrm{loc}}$, at the centre of the sheet. This can be converted to an effective mean strain, $\bar{\epsilon}_{\mathrm{eff}}$, using numerical results, see Fig. A4a. However, even without this conversion, the comparison between experiments and numerical results, see Fig. A4b is favourable.
@@
Extended Data Fig. A4: a, The strain measured locally at the middle of the sheet, $\epsilon_{\mathrm{loc}}$, over a longitudinal length of $l_{\mathrm{loc}}/W=0.25$ is smaller than the mean strain within the sheet, $\bar{\epsilon}$, but remains an approximately linear function of the strain. In all cases shown $L/W=2$ and $W_{\mathrm{load}}/W=0.25$. These parameters correspond to those used in the experiments. In the main text this relationship is used to convert the value of $\epsilon_{\mathrm{loc}}$ measured experimentally to an effective mean strain $\bar{\epsilon}_{\mathrm{eff}}$ for plotting in Fig. 3. Without this conversion, the comparison between experimental and numerical results is slightly worse as shown in b.
Mean strain versus imposed displacement
@@
Prior to buckling, the mean strain $\bar{\epsilon}=\Delta u_{y}/L$. However, after buckling, this relationship no longer holds. We therefore use numerical simulations to compute the relationship between the two: see Fig. A5a. In Fig. A5b we show see that the discrepancy between the two grows linearly above the buckling threshold, so that:
$\bar{\epsilon}-\frac{\Delta u_{y}}{L}=\begin{cases}0,\qquad\qquad\qquad\frac{\Delta u_{y}}{L}<\bar{\epsilon}_{c},\\ k\left(\frac{\Delta u_{y}}{L}-\bar{\epsilon}_{c}\right),\quad\frac{\Delta u_{y}}{L}>\bar{\epsilon}_{c}.\end{cases}$ (A1)
@@
Crucially, this means we can interchangeably use $\bar{\epsilon}-\bar{\epsilon}_{c}$ or $\Delta u_{y}/L-\bar{\epsilon}_{c}$ in scaling and experimental analyses. (Note, however, that the collapse of data is less satisfactory in the case $L=W$, for which the sheets are not particularly slender.)
@@
Extended Data Fig. A5: a, Mean strain, as measured from the arc-length, $\bar{\epsilon}$ is larger than the directly imposed strain $\Delta u_{y}/L$ in the post-buckling regime. b Up to the point of buckling $\bar{\epsilon}=\Delta u_{y}/L$, but even after buckling there is a linear relationship, as summarized in (A1). Here the data are scaled so that all data sets pass through the buckling point, $(\bar{\epsilon}_{c}t^{2}/W^{2},0)$; the inset shows the data before scaling.
Appendix D Details of Mathematical Modelling
@@
We consider an unstructured thin elastic sheet of thickness $t$, width $W$, and length $L$ ($t\ll W<L$) made of a material with Young’s modulus $E$ and Poisson’s ratio $\nu$. The sheet is clamped in a central segment along the short edges, of width $W_{\mathrm{load}}$, and subjected to an in-plane longitudinal stretching strain $\epsilon_{yy}$ — we choose the long edges to be parallel to the $y$-axis. Outside the clamps (i.e. for $|x|>W_{\mathrm{load}}/2$), the short edges are free, as well as the long edges (see Fig. A6a).
@@
$x$$y$$L$$W/2$$T$$T$bc
Extended Data Fig. A6: Schematic diagrams. a, A rectangular sheet, of width $W$ and length $L$, clamped a width $W_{\mathrm{load}}<W$ along its width. The sheet is longitudinally loaded at the clamps. b, A rectangular sheet, of width $W/2$ and length $L$, with localized load $T$ applied longitudinally at the corners $(0,\pm L/2)$. The sheet is transversely clamped at $(0,y)$. c, A cross-sectional profile (at $y=0$) of the sheet stretched along the $y$-axis by a localized load beyond buckling. The origin in the $\zeta$-axis corresponds to the reference plane of the planar solution.
We adopt the Föppl-von-Kármán (FvK) plate theory to model our sheet. We start by introducing the physical quantities that we will use. According to the FvK framework, we linearize the in-plane displacement-strain relationship:
@@
$\epsilon_{xy}=\frac{1}{2}\left(u_{y,x}+u_{x,y}+\zeta_{,x}\zeta_{,y}\right),$ (A2)
where $u_{x}$, $u_{y}$ are in-plane components of the displacement field, whose out-of-plane component is $\zeta$ (we use the comma notation in the subindex to denote the derivatives). We also invoke the two-dimensional Hookean constitutive relations, which link the in-plane strain to the in-plane stress ($\sigma_{\alpha\beta}$, which denotes the stress averaged across the sheet thickness):
@@
$\sigma_{xy}=\frac{1}{1+\nu}Y\,\epsilon_{xy},$ (A3)
@@
We consider the limit of small $W_{\mathrm{load}}$, $W_{\mathrm{load}}/W\ll 1$. Exploiting the symmetry along the $x$-axis, that we align with the short edge, we consider one half of this sheet, i.e. a flat sheet of width $W/2$ and length $L$ (see Fig. A6b). We examine next the instability of this sheet through a near-threshold analysis Xin and Davidovitch (2021a) of the planar state (i.e. $\zeta=0$). For the beyond threshold analysis, whose main parameters are represented in Fig. A6c, see the main text.
D.1 The planar state
@@
$\nabla^{4}\chi(x,y)=0,$ (A4)
after we eliminate the in-plane components of the stress tensor in favour of a single scalar function, the Airy potential $\chi(x,y)$, defined as
$\sigma_{xx}=\chi_{,yy};\,\sigma_{yy}=\chi_{,xx};\,\sigma_{xy}=-\chi_{,xy}.$ (A5)
We consider only small strains, and hence we assume the boundary conditions (BCs) to apply at the edges $x=0,W/2$ and $y=\pm L/2$. For convenience, we set the length $L=2$ and rescale $x,y$ with $L/2$. The set of BCs for the stress and the displacement fields is then:
at $y=\pm 1$: $\,\sigma_{yy}=\sigma_{xy}=0,$ (A6a)
at $x=W/2$: $\,\sigma_{xx}=\sigma_{xy}=0,$ (A6b)
at $x=0$: $\,u_{x}=0,\,u_{y,y}=\epsilon_{\text{planar}},$ (A6c)
which describes a sheet whose edge at $x=0$ is both transversely clamped ($u_{x}(x=0)=0$) and longitudinally strained ($\epsilon_{yy}(x=0)=u_{y}/L\equiv\epsilon_{\text{planar}}>0$). The edges at $x=W/2$ and $y=\pm 1$ are free.
We scale the stresses by the tension $T=\epsilon_{\text{planar}}Y$, $\sigma_{\alpha\beta}/T$, and we convert the set of BCs (A6) into BCs for the Airy potential invoking Eq. (A5):
at $y=\pm 1$: $\quad\chi=0\quad\text{and}\quad\chi_{,y}=0,$ (A7a)
at $x=W/2$: $\quad\chi=0\quad\text{and}\quad\chi_{,x}=0,$ (A7b)
at $x=0$: $\quad\chi_{,xx}-\chi_{,yy}\nu=1\quad\text{and}\quad\chi_{,xxx}+(2+\nu)\chi_{,xyy}=0.$ (A7c)
Note that we do not introduce any different notation for the scaled stresses $\sigma_{\alpha\beta}$, nor for the corresponding Airy potential $\chi$. Both equations in (A7a) and (A7b) follow from choosing the natural gauge of vanishing constants of integration when we integrate respectively the first equation in (A5) at $y=\pm 1$ and the second at $x=W/2$. This gauge is consistent with the BCs for the off-diagonal stress components at the free edges $y=\pm 1$ and $x=W/2$.
@@
We solve the bi-harmonic equation (A4) for the Airy potential with the set of BCs (A7) following the calculation of Benthem (1963) (further details on the method are given in Sec. D.2):
$\chi(x,y)=\text{Re}\Big{[}\sum_{k}\Big{\{}{\displaystyle C}_{k}\,e^{p_{k}(x-W/2)}+\hat{\displaystyle C}_{k}\,e^{-p_{k}x}\Big{\}}\Big{\{}\cos\left(p_{k}y\right)-(\cot p_{k})y\,\sin\left(p_{k}y\right)\Big{\}}\Big{]}\,,$ (A8)
@@
where ${\displaystyle C}_{k}$, $\hat{\displaystyle C}_{k}$ are constants given in (A27) in Sec. D.2, and $p_{k}$ are the (complex) roots of the trigonometric equation
$\sin 2p_{k}+2p_{k}=0,$ (A9)
@@
ordered by increasing $|p_{k}|$. We derive the Airy potential and we obtain the transverse and longitudinal components of the stress tensor (A5):
$\sigma_{xx}=\chi_{,yy}=$ $\text{Re}\Big{[}\sum_{k}\Big{\{}{\displaystyle C}_{k}\,e^{p_{k}(x-W/2)}+\hat{\displaystyle C}_{k}\,e^{-p_{k}x}\Big{\}}\Big{\{}-\left(p_{k}^{2}+2p_{k}\cot p_{k}\right)\cos\left(p_{k}y\right)+\cot p_{k}\,y\,\sin\left(p_{k}y\right)\Big{\}}\Big{]}\,,$ (A10a)
$\sigma_{yy}=\chi_{,xx}=$ $\text{Re}\Big{[}\sum_{k}\left(p_{k}\right)^{2}\Big{\{}{\displaystyle C}_{k}\,e^{p_{k}(x-W/2)}+\hat{\displaystyle C}_{k}\,e^{-p_{k}x}\Big{\}}\Big{\{}\cos\left(p_{k}y\right)-(\cot p_{k})y\,\sin\left(p_{k}y\right)\Big{\}}\Big{]}\,.$ (A10b)

### v2

@@ Appendix A Experimental methods
The stretching equipment is set up by two mechanical slider modules with $300\mathrm{~{}mm}$ range of motion. They are fixed to a stand and a clamping fixture is attached to each module (Fig. S1a). Both the fixtures and the stand were 3D printed (Raise3D FDM Printer) using polylactic acid (PLA). The specimen fixtures are customly designed to fit the slider and the clamp surfaces have teeth-like patterns to enhance grip (Fig. S1a, inset). The specimens (Fig. S1b) are made of laser-cut commercial polyvinyl chloride (PVC) sheets.
The elastic modulus of the sheets are determined from the tensile tests of rectangular strips with $W_{\mathrm{load}}=W$. The strips have, approximately, dimensions $W=20~{}\mathrm{mm},L=170~{}\mathrm{mm}$. Averaging five samples for each thickness, we obtain the Young’s modulus of PVC sheets with thickness $0.04,0.10,0.15\mathrm{~{}mm}$ to be $3.84,3.50,1.96\mathrm{~{}GPa}$, respectively (Fig. S1c). Note also that the value of the Young’s modulus only scales the forces required to cause a given deformation, not the deformation itself. The Poisson’s ratio of PVC sheets is 0.38.
@@
Extended Data Fig. S1: Experimental setups. a Tensile-stretching equipment consisting of linear-stage motors, 3D-printed specimen clamp fixtures and a stand (all labelled on the figure). b Testing specimen, showing the location of the strain gauge, cannula and where the sheet is clamped (see labels). c Stress-strain curves obtained from PVC sheets with fully clamped loading edge, $W_{\mathrm{load}}=W$; since these tests are performed in a tensile testing machine, the strain $\epsilon$ corresponds to $\epsilon_{0}$ used in the main text. d Composite image showing the evolution of the folding angle, $\alpha$, with increasing strain. e Visualization of the transverse curvature upon stretching via a projected laser line. Schematics illustrate the projection setup. The dotted line marks the straight line from which the beam deviates due to the curvature. f Visualization of the longitudinal curvature via a projected laser line.
The sheets used in the experiment have extended ‘tabs’ which are regions for clamping with $W_{\mathrm{load}}=10\mathrm{~{}mm}$ and $L_{\mathrm{load}}=30\mathrm{~{}mm}$ (Fig. S1b). The connection between the clamping regions and the main part of the sheets is designed with rounded corners of radius $1\mathrm{~{}mm}$ to prevent failure due to stress concentration. Tensile strains in the elastic sheets are measured by strain gauges (Tokyo Measuring Instruments Laboratory, GFLA-6-350-70) with gauge length of $6\mathrm{~{}mm}$, resistance of $350~{}\Omega$, and resolution of $1\mathrm{~{}\mu\epsilon}$. The strain gauges are attached to the centre of the specimen along the loading direction using an adhesive, and the gauge leads were connected to electrical wires that input to the strain data logger (Tokyo Measuring Instruments Laboratory, TDS-303). When clamping specimens between fixtures, vertical alignment is ensured for symmetric loading. Cannula needles with an outer diameter of $700\mathrm{~{}\mu m}$ are adhered to the $y$-middle of the sheet near the edges as an indicator in the measurement of inclination angles. Top-view photographs are taken of the specimens during the loading process (Fig. S1d), from which folding angles are measured using MATLAB. The doubly-curved shape of the sheet upon longitudinal stretching is shown by the curved projected laser lines along the two midlines of the sheet (Figs S1e,f), where two orthogonal laser beams are projected from a laser ruler placed above and to the side of a loaded sheet, as shown by the schematics.
Appendix B Experiments on crumpled paper
Experimental runs for three different crumpled sheets of paper are conducted using the same apparatus as for the plastic sheets. Sheets of printer paper ($80\mathrm{~{}gsm}$) of dimensions $W=8\mathrm{~{}cm}$, $L=16\mathrm{~{}cm}$ are first firmly crumpled by hand and are then loaded uniaxially with $W_{\mathrm{load}}=1.6\mathrm{~{}cm}$.
The angle $\alpha$ is measured directly as a function of the imposed longitudinal displacement, $\Delta u_{y}$ (see Eq. (1) in the main text), by measuring the distances between two markers with the initial distance of about $14\mathrm{~{}cm}$. This is then converted to the planar strain $\epsilon_{0}$, as in Eq. (2) of the main text. (It is not possible to measure strains locally using the strain gauge because the creases introduced by crumpling mean that the strain gauge cannot be glued tightly to the paper.) Inclination angles are measured in the same way as for plastic sheets—from needles attached at the middle of the long edges (refer to Fig. S1d).
Results of these experiment on crumpled paper are shown in Fig. S2 for three repeats (each on a fresh sheet of paper). These data show the same qualitative trend as experiments on uniform elastic sheets presented in the main text; however, there is more scatter in the data, which we attribute to the manual, and hence less repeatable, way in which crumpling is introduced in the experiment. (Since it is unclear what the corresponding stretching and bending moduli are for each sheet, we present the data only in their raw form here; the raw results of experiments on plastic sheets are reproduced with increased transparency to facilitate comparison.)
Extended Data Fig. S2: Experimental data for crumpled paper sheets. Results of tests on three crumpled-paper specimens (each with $L/W=2$, $W_{\mathrm{load}}/W=0.2$) as circles with different fills to distinguish each run. In this case, the measured strain is the planar strain $\epsilon_{0}$. Nevertheless, the observed folding angle generally displays a similar trend to that seen in experiments with plastic sheets (main text). The experimental data for plastic sheets is shown as the underlying transparent markers on the same plot for comparison, though we note that the strain measure for these data is actually $\epsilon_{\mathrm{loc}}$.
Appendix C Finite element method simulation
@@
Numerical simulations are conducted using the commercial finite element analysis software ABAQUS/2023. The model is constructed in 3D space using 2D shell sections and S4R elements. Linear elastic constitutive relation is applied with Young’s modulus $E=3000\mathrm{~{}MPa}$ and with Poisson’s ratio $\nu\in\{0.01,0.3,0.49\}$. Two analysis steps are conducted in sequence: first, a small out-of-plane displacement, $\Delta u_{z}$, is applied to the midline at $x=0$ of a magnitude of $\sim$ 1% of sheet thickness, creating initial perturbation to the planar sheet, while keeping both clamping edges fixed in all degrees of freedom; second, static analysis step is performed for the process of tensile-stretching induced folding of the elastic sheet. Displacement loading $\Delta u_{y}$ is applied to the bottom edge of the lower clamping region, while the top edge of the upper clamping region was fixed in all degrees of freedom.
@@
Extended Data Fig. S3: FEM simulation results for the effect of loading width on stress distribution. a Von-Mises stress distributions in sheets under localized tension, shown at $\bar{\epsilon}=\mathrm{0.005,0.012,0.025}$. Sheet geometry is $W_{\mathrm{load}}/W=0.1,L/W=2,t/W=2.5\times 10^{-3}$. b Out-of-plane deflections, $\zeta$, normalized by thickness, $t$, for sheets under localized loading (left), $W_{\mathrm{load}}/W=0.2$, and full-width loading (right), $W_{\mathrm{load}}/W=1$. c Transverse stress, $\sigma_{xx}$, in the same simulations as b correspondingly. b and c have $L/W=2,t/W=5\times 10^{-3}$, and are shown at the critical buckling strain for the localized loading case, i.e. $\bar{\epsilon}_{c}\approx 1.2\times 10^{-3}$.
@@
Extended Data Fig. S4: The scaled folding angle exhibits an intersection with the $x$-axis prior to the linear region, from where we determine the critical buckling strains $\bar{\epsilon}_{c}$ for FEM simulations. Results from FEM simulations are presented with a variety of values of $L/W$ and $W_{\mathrm{load}}/W$ (given in the legends of Fig. 2 in the main text), as well as $t/W$ (indicated by the color bar).
Appendix D Comparison of different strains in experiments
@@
Our numerical simulations allow us to compare the various types of strain that are measured experimentally.
@@
Firstly, our experiments measured the local strain, $\epsilon_{\mathrm{loc}}$, at the centre of the sheet. This can be converted to an effective mean strain, $\bar{\epsilon}_{\mathrm{eff}}$, using numerical results, see Fig. S5a. However, even without this conversion, the comparison between experiments and numerical results, see Fig. S5b, is favourable.
@@
Extended Data Fig. S5: a Numerical simulation results of the local strain at the centre of the sheet, $\epsilon_{\mathrm{loc}}$; to mimic the finite size of the strain gauge used experimentally, the strain is averaged over a longitudinal length $l_{\mathrm{loc}}/W=0.25$. Note that $\epsilon_{\mathrm{loc}}$ is smaller than the mean strain within the sheet, $\bar{\epsilon}$, but remains an approximately linear function of $\bar{\epsilon}$ throughout. In all cases shown $W_{\mathrm{load}}/W=0.25$, corresponding to the parameter values used in the experiments. In the main text this relationship is used to convert the value of $\epsilon_{\mathrm{loc}}$ measured experimentally to an effective mean strain $\bar{\epsilon}_{\mathrm{eff}}$. Without this conversion, the comparison between experimental and numerical results is slightly worse, as shown in panel b. Strain in the $x$-axis is $\bar{\epsilon}$ for simulation data points, but is $\epsilon_{\mathrm{loc}}$ for experiment data points (as originally measured). The dashed line shows the expected power-law $\alpha\propto\bar{\epsilon}^{3/4}$ for $\bar{\epsilon}\gg\bar{\epsilon}_{c}$; the solid curve shows the scaling law modified to accommodate the critical behavior of the system, i.e. Eq. (10) of the main text.
Mean strain versus imposed longitudinal displacement
@@
Prior to buckling, the mean strain along the centre line $\bar{\epsilon}=\epsilon_{0}=\Delta u_{y}/L$. However, after buckling, this relationship no longer holds. We therefore use numerical simulations to compute the relationship between the two: see Fig. S6a. In Fig. S6b we show that the discrepancy between the two grows linearly above the buckling threshold, so that:
$\bar{\epsilon}-\epsilon_{0}=\begin{cases}0,\quad\qquad\qquad\epsilon_{0}<\bar{\epsilon}_{c},\\ k\left(\epsilon_{0}-\bar{\epsilon}_{c}\right),\quad\epsilon_{0}>\bar{\epsilon}_{c}.\end{cases}$ (S1)
@@
Crucially, this means we can use $\bar{\epsilon}-\bar{\epsilon}_{c}$ or $\epsilon_{0}-\bar{\epsilon}_{c}$ interchangeably in scaling and experimental analyses. (Note, however, that the collapse of data is less satisfactory in the case $L=W$, for which the sheets are not particularly slender.)
@@
Extended Data Fig. S6: a Mean strain measured from the arclength, $\bar{\epsilon}$, is larger than the directly imposed strain, ϵ0subscriptitalic-$\epsilon_{0}\equiv$ $\Delta u_{y}/L$, in the post-buckling regime. b Up to the point of buckling $\bar{\epsilon}=\epsilon_{0}$, but even after buckling there is a linear relationship $\bar{\epsilon}-\epsilon_{0}\propto\epsilon_{0}-\bar{\epsilon}_{c}$, as summarized in (S1). Here the data are scaled so that all data sets pass through the buckling point, $(\bar{\epsilon}_{c}W^{2}/t^{2},0)$; the inset shows the data before scaling.
Extended Data Fig. S7: Numerical simulation results for the inclination angle with different Poisson’s ratios. a Raw results showing the inclination angle as a function of strain with different thickness ratios (see legends). b Rescaling the imposed strain by the Poisson’s ratio-dependence in the critical strain provides some collapse of data with different $\nu$. c Better collapse of the data is obtained by rescaling the $y$-axis as expected based on our theory. The dashed line shows the expected power-law $\alpha\propto\bar{\epsilon}^{3/4}$ for $\bar{\epsilon}\gg\bar{\epsilon}_{c}$; the solid curve shows the scaling law modified to accommodate the critical behavior of the system, i.e. Eq. (10) of the main text.
Appendix E Details of Mathematical Modelling
@@
We consider an unstructured thin elastic sheet of thickness $t$, width $W$, and length $L$ ($t\ll W<L$) made of a material with Young’s modulus $E$ and Poisson’s ratio $\nu$. The sheet is clamped in a central segment along the short edges, of width $W_{\mathrm{load}}$, and subjected to an in-plane longitudinal stretching strain $\epsilon_{yy}$ — we choose the long edges to be parallel to the $y$-axis. Outside the clamps (i.e. for $|x|>W_{\mathrm{load}}/2$), the short edges are free, as are the long edges (see Fig. S8a).
@@
[…]
Extended Data Fig. S8: Schematic diagrams. a A rectangular sheet, of width $W$ and length $L$, clamped a width $W_{\mathrm{load}}<W$ along its width. The sheet is longitudinally loaded at the clamps. b A rectangular sheet, of width $W/2$ and length $L$, with localized load $T_{0}$ applied longitudinally at the corners $(0,\pm L/2)$. The sheet is transversely clamped at $(0,y)$. c A cross-sectional profile (at $y=0$) of the sheet stretched along the $y$-axis by a localized load beyond buckling. The origin in the $\zeta$-axis corresponds to the reference plane of the planar solution.
We adopt the Föppl-von-Kármán (FvK) plate theory to model our sheet. We start by introducing the physical quantities that we will use. According to the FvK framework, we use the in-plane displacement-strain relationship:
@@
$\epsilon_{xy}=\frac{1}{2}\left(u_{y,x}+u_{x,y}+\zeta_{,x}\zeta_{,y}\right),$ (S2)
where $u_{x}$, $u_{y}$ are in-plane components of the displacement field, whose out-of-plane component is $\zeta$ (we use the comma notation in the subindex to denote the derivatives). We also invoke the two-dimensional Hookean constitutive relation, which links the in-plane strain to the in-plane stress ($\sigma_{\alpha\beta}$, which denotes the stress averaged across the sheet thickness and where indices $\alpha,\beta$ can be either $x$ or $y$):
@@
$\sigma_{xy}=\frac{1}{1+\nu}Y\,\epsilon_{xy},$ (S3)
@@
We consider the limit of small $W_{\mathrm{load}}$, $W_{\mathrm{load}}/W\ll 1$. Exploiting the symmetry along the $x$-axis, that we align with the short edge, we consider one half of this sheet, i.e. a flat sheet of width $W/2$ and length $L$ (see Fig. S8b). We examine next the instability of this sheet through a near-threshold analysis Xin and Davidovitch (2021a) of the planar state (i.e. $\zeta=0$). For the beyond threshold analysis, whose main parameters are represented in Fig. S8c, see the main text.
The planar state
@@
$\nabla^{4}\chi(x,y)=0,$ (S4)
after we eliminate the in-plane components of the stress tensor in favour of a single scalar function, the Airy potential $\chi(x,y)$, defined such that
$\sigma_{xx}=\chi_{,yy};\,\sigma_{yy}=\chi_{,xx};\,\sigma_{xy}=-\chi_{,xy}.$ (S5)
We consider only small strains, and hence we assume the boundary conditions (BCs) to apply at the edges $x=0,W/2$ and $y=\pm L/2$. For convenience, we rescale $x,y$ by $L/2$ so that the edges are at $y=\pm 1$. The set of BCs for the stress and the displacement fields is then:
at $y=\pm 1$, $x\neq 0$: $\,\sigma_{yy}=\sigma_{xy}=0,$ (S6a)
at $x=W/2$: $\,\sigma_{xx}=\sigma_{xy}=0,$ (S6b)
at $x=0$: $\,u_{x}=0,\,u_{y,y}=\epsilon_{0},$ (S6c)
which describes a sheet whose edge at $x=0$ is both transversely clamped ($u_{x}(x=0)=0$) and longitudinally strained ($\epsilon_{yy}(x=0)=\Delta u_{y}/L\equiv\epsilon_{0}>0$, with $\Delta u_{y}\equiv u_{y}(x=0,y=L/2)-u_{y}(x=0,y=-L/2)$). The edges at $x=W/2$ and $y=\pm 1$ are free.
We scale the stresses by the characteristic tension $T_{0}=Y\epsilon_{0}$ (with $Y=Et$), and so let $\sigma_{\alpha\beta}/T_{0}$ be the dimensionless stress tensor. Converting the set of BCs (S6) into BCs for the Airy potential, via (S5), we find:
at $y=\pm 1$, $x\neq 0$: $\quad\chi=0\quad\text{and}\quad\chi_{,y}=0,$ (S7a)
at $x=W/2$: $\quad\chi=0\quad\text{and}\quad\chi_{,x}=0,$ (S7b)
at $x=0$: $\quad\chi_{,xx}-\nu\chi_{,yy}=1\quad\text{and}\quad\chi_{,xxx}+(2+\nu)\chi_{,xyy}=0.$ (S7c)
(Note that we do not introduce any different notation for the dimensionless stresses $\sigma_{\alpha\beta}$, nor for the corresponding Airy potential $\chi$, but work in dimensionless terms henceforth.) Both equations in (S7a) and (S7b) follow from choosing the natural gauge of vanishing constants of integration when we integrate respectively the first equation in (S5) at $y=\pm 1$ and the second at $x=W/2$. This gauge is consistent with the BCs for the off-diagonal stress components at the free edges $y=\pm 1$ and $x=W/2$.
@@
We solve the bi-harmonic equation (S4) for the Airy potential with the set of BCs (S7) following the calculation of Benthem Benthem (1963) (further details on the method are given in the next subsection). In short, the general solution may be written
$\chi(x,y)=\text{Re}\Big{[}\sum_{k}\Big{\{}{\displaystyle C}_{k}\,e^{p_{k}(x-W/2)}+\hat{\displaystyle C}_{k}\,e^{-p_{k}x}\Big{\}}\Big{\{}\cos\left(p_{k}y\right)-(\cot p_{k})y\,\sin\left(p_{k}y\right)\Big{\}}\Big{]}\,,$ (S8)
@@
where ${\displaystyle C}_{k}$, $\hat{\displaystyle C}_{k}$ are constants given in (S27) in Sec. E, and $p_{k}$ are the (complex) roots of the trigonometric equation
$\sin 2p_{k}+2p_{k}=0,$ (S9)
@@
ordered by increasing $|p_{k}|$. Differentiating the Airy potential, we obtain the transverse and longitudinal components of the stress tensor (S5):
$\sigma_{xx}=\chi_{,yy}=$ $\text{Re}\Big{[}\sum_{k}\Big{\{}{\displaystyle C}_{k}\,e^{p_{k}(x-W/2)}+\hat{\displaystyle C}_{k}\,e^{-p_{k}x}\Big{\}}\Big{\{}-\left(p_{k}^{2}+2p_{k}\cot p_{k}\right)\cos\left(p_{k}y\right)+\cot p_{k}\,y\,\sin\left(p_{k}y\right)\Big{\}}\Big{]}\,,$ (S10a)
$\sigma_{yy}=\chi_{,xx}=$ $\text{Re}\Big{[}\sum_{k}\left(p_{k}\right)^{2}\Big{\{}{\displaystyle C}_{k}\,e^{p_{k}(x-W/2)}+\hat{\displaystyle C}_{k}\,e^{-p_{k}x}\Big{\}}\Big{\{}\cos\left(p_{k}y\right)-(\cot p_{k})y\,\sin\left(p_{k}y\right)\Big{\}}\Big{]}\,.$ (S10b)